Shift summary for Evening of July 14th

Leo Bellantoni (bellanto@fnpxsr)
Thu, 15 Jul 1999 00:05:12 CDT

Shift Summary for Eve shift, 14/Jul/99

On Shift:
Leo Bellantoni, Collin Wolfe

Hours of beam delivered at all intensities: 6.5
Hours of beam delivered at nominal intensity: 6.5
Hours of E832 data (special runs not included): 6.4
Hours of special runs (laser scans, muon runs...): 0.3

Accounting of lost time:
TeVatron down with concurrent monitoring / DAQ problems
Switching from running laser scan to data taking.

========================================================================
Runs ended this shift:
Run 13907 - 21.6M events - BEAM_832

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift:
None

========================================================================
Overview of shift:
E832 datataking, banana maps upgraded.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13907 120 15:30 Beam expected in 10 minutes.
129 15:40 It's back!
163 Gone.
188 17:01 Back.
200 CsI channel 961 has accumulated 42 comparator
errors at this point.
228 18:00 SEM wanders over 7e12; we phone MCR.
near 21:10 Suspect that monitor histograms are not
being updated...
21:25 Beam disappears
308 21:28 Indeed, an L3 process has died on Ktev3. Run
stopped.

----- near 21:30 opcrc error on start; must drt_fresh_start
Beam disappearance attributed to TeV power
supply problem. No estimate.

21:48 Up and running again. Us, not the TeVatron.
Bananas modified: smaller intime window and
new kumacs.

13910 1 22:43 Laser scan starts
22:58 TeVatron back. Turns out it was just a coupla
blown fuses. We abort the laser scan.

13911 1 23:08 Running again in E832 mode.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Laser scan if down more than 1 hr
No polarity flip or muon run at midnight
Red SDAQ buttons "RR-160-NIM2" and "UCLA-PS4" may be ignored